OpenStudy (anonymous):

\[\sum_{0}^{\infty} 1\div \left( \sqrt{2} \right)^{n}\]

OpenStudy (anonymous):

did u check if it converges?

OpenStudy (anonymous):

no how do i go about that

OpenStudy (anonymous):

It's a geometric series, since sqrt(2) > 1, it converges, to a formula you really should look up and memorize.

OpenStudy (anonymous):

the one where it says -1<r<1?

OpenStudy (anonymous):

and since it is convergent how do i find its sum?

OpenStudy (anonymous):

No, I'll give it: \[\sum_{n=0}^{\ \infty}(\frac{1}{k})^n=\frac{1}{1-\frac{1}{k}}\]

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i thought it was more complicated

OpenStudy (anonymous):

Nope. Pretty straight-forward.

OpenStudy (anonymous):

wow thank you

OpenStudy (anonymous):

where is the good answer button haha?

OpenStudy (anonymous):

Right side of the tan bar with your name in it.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Any time.